A lift with open top is moving up with acceleration f. A body is thrown up with speed u relative to lift and it comes back to lift in time t. Then speed u is ...
{g = acceleration due to gravity}

Grade:

1 Answers

kumar kanishk
6 years ago
in a lift net acceleration = a + g
                                      = f + g
time of flight t = 2u / f+g
                     ft + gt = 2u
             so
                  u = ft/2 + gt/2